sin z 21) 20) Cosx dx (r? + 1) X 22) sin mx dx 2(x² + a²² (a > 0, b>0) 23) cos ex - cos bx -dx x?

use residue theorem to evaluate the following integrals
Answer #1

use residue Theoren to evaluate DATE The following integral 2 202 CosX dx 2 o (21) 2 solchon Consider the integral 17 2 iz aDATE re ob ud denotes Poles withine Where ERT the residue Plz) Which the lig etc (zn)? whicet Now Take 4 127 e a pole of withod s 2 ina a e ooda = 20 i Ent (27) =znice) att i se 00 2 in 2 das T l ze In 2 2 2 cavating real con marion part on both sideDATE -R to-r (1) to real any (۱) The upper hall of the circl 12128 Show in figure below X CR R r Zoo CA rR since f(a) has noROB 2 2 2 2 0 25 ZRO * do Colo</2 con gyott leh) T which cy R 49 Also I field2 co Petre copot things 4 do ié et og ryo 2 HenDATE 3 Sirona da 22 اما را 0 а (-2, y2 solution consider the integral zlezza f(e) dz elona where f(2)= the given functiva hayI of Cees Ti ri enam ? au lmon or e 00 2a3 002 I sonen dy Xina 22 Ti ema (me auating imaginary on both sedes Joo sim ma ( meDATE (oo) I todarf flugdrt / fonde f(2) dra fu = f (2) de sino V 7374 +1z) 0 = -21 i Lé (mondes 4.3 Seime 2/27aza ht =0 2(2²